# A generic transformation is invertible

### v3

@@ 1. Introduction
Let $(X,\mu)$ be a standard non-atomic probability space and $T:X\to X$ be measure preserving, i.e., satisfies $\mu(T^{-1}(A))=\mu(A)$ for every measurable $A\subset X$. Recall that the corresponding Koopman operator [^1] We denote it by the same letter as the transformation following a tradition in ergodic theory. $T$ on $L^{2}(X,\mu)$ is defined by $(Tf)(x):=f(Tx)$ and is an isometry on $L^{2}(X,\mu)$. Moreover, this operator is invertible, i.e., unitary, if and only if222see, e.g., [7, Rem. 8.48(d)](e) for the “only if” direction the underlying transformation is essentially invertible.
@@
In particular, the set $\mathcal{T}_{\text{inv}}$ is residual in $\mathcal{T}$. After the proof of the dense $G_{\delta}$ property in Section 2 we also present an alternative, more elementary and direct proof of residuality in Section 3.
@@
Acknowledgement. The author is indebted to Valentin Gillet for very helpful discussions regarding the points of continuity method in the operator theoretic context and, in particular, for the references [9, 10].
2. Proof
@@
By a classical result of Baire, see Kechris [14, Thm. 24.14], the set of continuity points of a Baire class $1$ map [^7] A map $f:X\to Y$ between topological spaces $X$ and $Y$ is called of Baire class $1$ (or Borel $1$) if the preimage of every open subset of $Y$ is $F_{\sigma}$ in $X$. $f:X\to Y$ form a dense $G_{\delta}$ subset of $X$ whenever $X,Y$ are metrizable and $Y$ is separable. By Lemmata 2.2 and 2.3 it just remains to check the Baire class $1$ property of the identity map $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$. This can be done as in [10, Lemma 2.9] and we repeat here the simple argument for completeness.
@@
3. Alternative proof of residuality
We now present an alternative, more elementary and direct proof of the residuality of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ using Lemma 2.1 and argue as in [6]. Let $(f_{j})$ be dense in $L^{2}(X,\mu)$. Since every isometry has closed range,
(2) $\mathcal{T}\setminus\mathcal{T}_{\text{inv}}=\bigcup_{j,k\in\mathbb{N}}\left\{T\in\mathcal{T}:\,\text{dist}(f_{j},\text{rg}T)>\frac{1}{k}\right\}=:\bigcup_{j,k\in\mathbb{N}}M_{j,k}.$
We will show that each set $M_{j,k}$ is nowhere dense. By density of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ it is enough to show $\overline{M_{j,k}}\cap\mathcal{T}_{\text{inv}}=\emptyset$. Assume the opposite and let $T\in\mathcal{T}_{\text{inv}}$ satisfy $\lim_{n\to\infty}T_{n}=T$ strongly for some $(T_{n})\subset\mathcal{T}$ with $\text{dist}(f_{j},\text{rg}T_{n})>\frac{1}{k}$ for every $n\in\mathbb{N}$. Then $g:=T^{-1}f_{j}$ satisfies
$\lim_{n\to\infty}T_{n}g=Tg=f_{j},$
i.e., $\lim_{n\to\infty}\text{dist}(f_{j},\text{rg}T_{n})=0$, a contradiction.

### v4

@@ 1. Introduction
Let $(X,\mu)$ be a standard non-atomic probability space and $T:X\to X$ be measure preserving, i.e., satisfies $\mu(T^{-1}(A))=\mu(A)$ for every measurable $A\subset X$. Recall that the corresponding Koopman operator [^1] We denote it by the same letter as the transformation following a tradition in ergodic theory. $T$ on $L^{2}(X,\mu)$ is defined by $(Tf)(x):=f(Tx)$ and is an isometry on $L^{2}(X,\mu)$. Moreover, this operator is invertible, i.e., unitary, if and only if222see, e.g., [7, Rem. 8.48(e)] for the “only if” direction the underlying transformation is essentially invertible.
@@
In particular, the set $\mathcal{T}_{\text{inv}}$ is residual in $\mathcal{T}$. We present two proofs of the $G_{\delta}$-property, a short one in Section 3 based on the fact that an isometry is invertible if and only if it has dense range, and a longer one in Section 4 using the points of continuity method.
@@
Acknowledgements. The author is very grateful to Ethan Akin for a crystallisation of the argument in Section 3 from its original version concerning residuality. She is also indebted to Valentin Gillet for very helpful discussions regarding the points of continuity method in the operator theoretic context and, in particular, for the references [9, 10].
2. Proof: Density
@@
3. First Proof of $G_{\delta}$ Property
The first proof of the $G_{\delta}$ property of $\mathcal{T}_{\text{inv}}$ in $\mathcal{T}$ is inspired by [6].
Let $(f_{j})$ be dense in $L^{2}(X,\mu)$. Since every isometry has closed range and $(Tf_{j})$ is dense in it, we have
$\mathcal{T}_{\text{inv}}=\bigcap_{j,k\in\mathbb{N}}\bigcup_{l\in\mathbb{N}}\left\{T\in\mathcal{T}:\,\|f_{j}-Tf_{l}\|<\frac{1}{k}\right\}=:\bigcap_{j,k\in\mathbb{N}}\bigcup_{l\in\mathbb{N}}M_{j,k,l}.$
It is clear that every set $M_{j,k,l}$ is open for the strong operator topology, finishing the proof.
4. Second Proof of $G_{\delta}$ Property
We also provide an alternative proof of the $G_{\delta}$ property via the points of continuity method.
@@ 4. Second Proof of $G_{\delta}$ Property
By a classical result of Baire, see Kechris [14, Thm. 24.14], the set of continuity points of a Baire class $1$ map [^7] A map $f:X\to Y$ between topological spaces $X$ and $Y$ is called of Baire class $1$ (or Borel $1$) if the preimage of every open subset of $Y$ is $F_{\sigma}$ in $X$. $f:X\to Y$ form a dense $G_{\delta}$ subset of $X$ whenever $X,Y$ are metrizable and $Y$ is separable. By Lemmata 4.1 and 4.2 it just remains to check the Baire class $1$ property of the identity map $\mathrm{id}:(\mathcal{T},\texttt{SOT})\to(\mathcal{T},\texttt{SOT}^{*})$. This can be done as in [10, Lemma 2.9] and we repeat here the simple argument for completeness.